# High-precision measurements of electric-dipole-transition amplitudes in excited states of 208Pb using Faraday rotation spectroscopy

### v1

@@ III.2 Data collection
For each E1/M1 comparison experiment, after the furnace reached equilibrium and laser beam overlap through the sample was optimized, we performed sequential frequency scans of the two lasers (both upward and downward going) across the respective transitions collecting $1f$ and $2f$ lock-in outputs and Fabry-Pérot transmisison data both with and without the DC Faraday magnetic field. The eight-scan sequence required roughly one minute to complete and was repeated (in various orders) over the course of roughly 45 minutes. In order to calibrate the measured $1f/2f$ lock-in ratio in terms of absolute radians (effectively measuring the constant $\kappa$ in Eq. 9), both before and after each such set of runs we induced a common angular ‘uncrossing’ in our polarimeter by using a computer-controlled stepper motor and differential micrometer to effect a small, controlled polarization rotation of roughly 5 mRad in our second polarizer (analyzer). This calibration procedure was performed twice before and twice after each data set. Since we measure the ratio of amplitudes, the exact magnitude of the mechanical uncrossing is not important, only that it is the same for both lasers. The precision and reproducibility of this procedure was aided by the microradian-level control of our mechanical polarizer mount (see Sec. IV.C.). As discussed below, we assigned calibration errors based on both the reproducibility of repeated calibration steps and the variance of the before vs. after data-set calibration exercises. For every scan we also collect Fabry-Pérot transmission spectra. We also record the value of the DC solenoid current applied, as well as the temperature readings for each of the three thermocouples within the interaction region.
We typically re-optimized the optical system after one data set. Over the course of several weeks of data collection for each comparison experiment, we changed a variety of experimental parameters such as the magnitude of the DC Faraday field, the E1 laser power, laser sweep conditions, etc. Data were taken over a wide range of sample temperatures between 700-$850\,^{\circ}$ which corresponds in both comparison experiments to an order-of-magnitude change in Pb vapor density, and a factor of four change in the thermal Boltzmann factor noted in the previous section. In the end, we collected and analyzed more than 500 individual spectral scan pairs for each comparison experiment taken over a wide variety of experiment conditions including a range of different DC magnetic fields and E1 laser powers. A comprehensive analysis of both statistical and potential systematic errors associated with the overall data sets is discussed below.
@@ IV.1 Line shape fitting procedure
Figure 4: Sample data (blue markers) for single (field on - field off) Faraday rotation scans for all three transitions at $\,{}^{\circ}$ and 44 Gauss. Each scan displayed required roughly 30 seconds to acquire. The line of best fit is shown in red. While the 1279 nm and 368 nm optical rotation amplitudes are of comparable size, the amplitude of the 406 nm transition is an order of magnitude smaller due to the reduced thermal population of $(6p^{2})^{3}P_{2}$ level compared with the $(6p^{2})^{3}P_{1}$ level.
@@
Figure 5: Sample single transmission data scans (blue markers) and line of best fit (red) at $\,{}^{\circ}$ for the 1279 nm and 368 nm transitions. The noticeably higher noise level for the 1279 nm transition is a consequence of the much smaller $2f$ modulation depth due to the reduced Verdet constant of the CeF3 crystal at 1279 nm.
@@
As a first step in preparing our spectra for analysis, a linearized and calibrated frequency axis of each spectra is obtained by fitting the Fabry-Pérot transmission data to an Airy distribution. Nonlinear scan behavior was accounted for by expanding the Airy argument up to a fourth-order polynomial. Higher-order polynomials were explored but were found to have no effect on the extracted E1 transition matrix element. The accurately-measured free spectral range (FSR) of the Fabry-Pérot cavities allowed absolute frequency-scale calibration of sufficient accuracy for our needs. Next, we subtract the lock-in $1f/2f$ signals with magnetic field off from the associated field-on spectra which removes frequency-dependent polarization backgrounds unrelated to the atoms. We then use the average value of the before and after data set calibration steps to provide background-subtracted spectra with amplitudes in units of radians. Using the calibrated frequency scale, we can separately create transmission spectra suitable for direct optical depth analysis using the $2f$ lock-in output for the case of zero magnetic field (that is, using our ‘background’ scans).
@@
We establish a nominal temperature to associate with each scan for two reasons. First, as described next, we are then able to fix the Doppler width for our eventual fits based on that temperature, simplifying our fitting procedure. Second, as shown in Eqs. 4 and 8, we require a temperature value to evaluate the Boltzmann factor that allows us to ultimately extract the E1 amplitude values. For the large majority of the data collected, the three thermocouple reading within the vapor cell enclosure were consistent within the quoted ±$\pm 1.5\,^{\circ}$ thermocouple accuracy. Here we assigned the central thermocouple value (labeled “B” in Fig. 3) to be the nominal temperature. We ultimately assigned an overall temperature-uncertainty systematic error to our results, as described in Sec. IV.C.
@@
Finally, we also establish a nominal DC magnetic field value to associate with each Faraday scan so that we can fix the Zeeman splitting in our fits as well. In Appendix B we describe our method to calibrate the magnetic field based on a Doppler-free, two-step spectroscopy scheme using the 1279 nm M1 laser overlapping with the 368 nm UV laser to probe the Zeeman structure of the $(6p^{2})^{3}P_{1}$ whose magnetic $g$-factor is very well known (see Wood1968 and references therein). Sec. IV.C. further discusses the accuracy of this method and potential systematic errors associated with possible inaccuracies in assigning magnetic field values to our fits.
@@
For the case of the direct transmission fits, we similarly fix the Doppler widths and then fit the (magnetic-field-off) $2f$ lock-in signals to the function described by Eqs. 1 and 2. We allow a frequency-dependent linear intensity background in the fits to account for incident laser power variation across the resonance. From these fits we again extracted homogeneous width components and also peak optical depths from which the E1 amplitudes can once again be obtained. Importantly, we saw good agreement of the homogeneous widths as extracted from the Faraday and the transmission line shapes. For experimental conditions where direct transmission analysis was statistically feasible, these quite distinct analysis procedures provide a powerful test of potential hidden systematic errors. For example, the transmission fits do not rely on overall amplitude or magnetic field calibration, whereas the Faraday technique provides a zero-background signal whose very high intrinsic precision allows data collection over a much wider range of temperatures and cell vapor densities. For the 368 nm transition, originating from the $(6p^{2})^{3}P_{1}$ state, we were able to obtain high-quality transmission fits and extract accurate E1 amplitude values for the data taken at $\,{}^{\circ}$ and $\,{}^{\circ}$. Because of the order-of-magnitude smaller thermal Boltzmann factor for the $(6p^{2})^{3}P_{2}$ state, we were able to obtain transmission results with precision sufficient for useful statistical comparison to the Faraday results only at the highest temperature ($\,{}^{\circ}$) for 406 nm E1 transition. Fig. 4 shows samples of single, background-subtracted Faraday rotation scans with solid lines representing model fits. These sample scans were taken with nominal values of magnetic field and laser power, and at a temperature in the middle of our experimental range. Each spectrum required roughly 30 seconds to acquire. Fig. 5 shows a sample of direct transmission spectra for the 368 nm E1/M1 comparison experiment taken at $\,{}^{\circ}$. Data points and model fit are again shown.
@@ IV.2 Summary of results
After amplitude values from the Faraday fits (i.e. the $\mathcal{C}$ constant in Eq. 7), we insert the nominal temperature for the particular spectrum pair into the Boltzmann factor expression in Eq. 8 along with the other constants, and the accurately-known Porsev2016 M1 reduced matrix element $\langle||M1||\rangle=1.293(1)$ a.u. to obtain a value for the E1 reduced matrix element. When we analyze a full set of scans within one data set (typically 25 up and 25 down spectrum pairs), the fractional statistical standard error for the average extracted E1 transition amplitude is at or below $1\%$. Similarly, from a pair of fitted transmission spectra, we insert the values of the optical depths into Eq. 4 along with the thermal Boltzmann factor to again obtain E1 transition amplitudes.
@@
In Fig. 6 we have assembled results of all our data sets with associated error bars grouped in order of increasing cell temperature/density for each of the E1 transitions including Faraday results (red circles) as well the transmission line shape results (blue triangles) completed for the highest temperatures in each case. Over an order of magnitude change in Pb vapor density, we find that there is no resolved trend in the E1 amplitude results versus temperature in either case, and that the transmission analysis is in excellent agreement with the Faraday results where common analysis was possible. Fig. 6 includes statistical errors only. For each data set we found the average of the set of individual scan weighted by the statistical errors derived from the fit procedure, and then added that data set fractional uncertainty in quadrature with the uncertainty in our mechanical angle calibration procedure (typically a few tenths of 1% fractional uncertainty, and common to all scans within a set). For this latter contribution we considered the difference between before-set and after-set calibration procedures. As indicated in Table 1, the final fractional statistical error for the complete data sets for each E1 transition amplitude was near 0.2%. We determined our final value for each of our two amplitudes by taking the weighted mean value of all data sets whose individual uncertainties were computed from the standard error arising from fitted values from the individual scans comprising that set. As an alternative method we measured the mean of the (unweighted) histogram consisting of every individual measurement of the transition amplitude. In the case of both transitions we found excellent agreement in the mean values derived from these two methods. Our final values including statistical errors only are: $\langle||E1||\rangle_{368}=1.897(4)$ a.u. and $\langle||E1||\rangle_{406}=3.008(7)$ a.u.
@@ IV.3 Systematic error discussion
In general we took two different approaches in exploring our results for potential systematic errors. First, we split the overall collected data into discrete subsets taken with differing experimental conditions and looked for evidence of statistically significant dependencies on various parameters. Examples of this include considering results from upward vs. downward laser scans, results for various values of DC magnetic fields used for Faraday rotation data, different E1 laser powers, and as suggested in Fig. 6, results grouped by cell temperature. Second, to explore possible line shape model or fitting inaccuracies, we refit subsets of data with altered model parameters or fitting procedures and looked for changes in the extracted values for our E1 matrix elements. Athough the large majority of the complete body of data for each comparison experiment were Faraday rotation-based, we nevertheless separated the analysis of the Faraday and transmission results to account for distinct systematic errors. While the Faraday results were far more statistically precise, they were also subject to more potential systematic errors. We here enumerate and summarize the systematic error contributions to the Faraday data only in Table 1. We note that the final two entries in the table (in red text) are factors which directly scale the extracted E1 amplitude (see Eqs. 4 and 8), and would potentially bias both Faraday and transmission data in the same direction. We therefore took the approach of analyzing statistical and all relevant systematic contributions for Faraday data and transmission data separately, took the weighted average of these, and then finally combined that resultant error with the contributions due to temperature and M1 amplitude uncertainty in quadrature to obtain the final quoted uncertainty discussed in Sec. V. Below we focus largely on the Faraday data systematics investigations in particular.
@@ IV.3.1 Line shape model and fits
Line shape fits were also performed both with and without the off-resonant ‘wings’ of the Faraday spectra, and while the observed changes were small, the effect on the 368 nm Faraday spectra was slightly larger, due to the complications associated with small ellipticity and asymmetry caused by the CeF3 crystal modulator Lacy2024 . For these 368 nm Faraday line shape fits, an additional ellipticity parameter was included to account for this imperfection which has the effect of producing a small difference in the amplitudes of the $\sigma_{+}$ and $\sigma_{-}$ contributions to the (otherwise symmetric) Faraday spectra. Typically, these effects were statistically unresolved, which was achieved by ensuring the lasers were aligned along the optic-c axis of the crystal (see Lacy2024 for further details). To explore possible correlations of this ellipticity with Faraday amplitude and eventually matrix element determination, we compared fits of a representative batch at $\,{}^{\circ}$ both with and without an ellipticity fit parameter and found changes at the 0.1% level. Accounting for such ellipticity was not required for the 406/1279 nm $\langle\norm{E1}\rangle$ comparison, since the biaxial TGG crystal modulator used there exhibits no birefringenece. Another potential feature in all of our Faraday line shapes is due to possible imperfect background subtraction which we explored by re-fitting subsets of data with a linear background slope parameter. The results of these studies are included in Table 1. We note that the first three entries in the table apply also to the analysis of possible systematic errors in our transmission spectra, and we took a similar approach to re-fitting these spectra, with error contributions at a similar 0.1% level or below.
@@ IV.3.2 Calibration uncertainties
In our Faraday rotation line shape analysis, we enter the particular Zeeman shifts in the component dispersion curves as fixed, ‘known’ parameters while we extract the line shape amplitude and homogeneous widths from the fits themselves. If either the magnetic field calibration or the values of the relevant $g$-factors that we input are incorrect, this could lead to systematic errors in key fitted parameters. The various $g$-factors discussed and referred to in Sec II have been measured Wood1968 ; Lurio1970 and are in good agreement with theory values Porsev2016 . We use the following consensus values: $g=1.501(2);\,g^{\prime}=1.276(2);\,g^{\prime\prime}=1.352(3)$. Residual errors in these values do not contribute to the 368/1279 nm comparison since they involve a common $g$-factor. To study the impact of these uncertainties on the 406/1279 nm comparison, using both experimental and simulated data, we refit the spectra with $g$-factors altered by an amount consistent with the error bars stated above, and assessed the impact of these uncertainties on the eventual computed value for the E1 amplitude (see Table 1). Regarding magnetic field calibration, our in situ method for field calibration at the 1% level of accuracy is described in Appendix B. We can similarly explore the consequence of incorrectly fixed magnetic field values through refitting experimental or simulated data with various slightly altered fixed values for the field. In the limit of the small-field approximation, the effect of a changing magnetic field would exactly cancel in our ratio measurement; here there is some differential effect since the Zeeman splitting represents a different fraction of the overall line shape width for the M1 vs E1 cases.
@@
The most significant uncertainty in our entire experimental project, equally relevant for Faraday and Transmission data analysis, is associated with temperature calibration and the effect that this has on the thermal Boltzmann factor appearing in Eqs. 4 and 8. As discussed in Sec. III, the sample temperature was monitored with three s-type thermocouples located in the stainless stell enclosure containing our vapor. These have quoted accuracy of $\Delta T=\pm$ $\,{}^{\circ}$. In most cases, the three thermocouple readings were in agreement within this intrinsic error. In these instances, the temperature uncertainty was $\Delta T/\sqrt{3}\approx 0.9$$\,{}^{\circ}$. For the 368 nm and 406 nm transitions, this uncertainty propagates to a Boltzmann factor uncertainty of 0.46% and 0.58% respectively at $\,{}^{\circ}$ where the bulk of the data were collected (this percentage error depends very weakly on temperature over our full experimental range). In the few cases where there was a small but resolvable temperature gradient across the atomic sample, we assigned an additional temperature uncertainty (up to $\pm$ $\,{}^{\circ}$) whose additional propagated uncertainty in the transition amplitude value was added in quadrature with fit errors and calibration errors, effectively reducing the statistical weight of that run.

### v2

@@ III.2 Data collection
For each E1/M1 comparison experiment, we performed sequential frequency scans of the two lasers (both upward and downward going) across the respective transitions both with and without the DC Faraday magnetic field. The eight-scan sequence required roughly one minute to complete and was repeated (in various orders) over the course of roughly 45 minutes. For each scan, we also collected Fabry-Pérot transmission spectra, and recorded the value of the applied DC solenoid current, as well as the temperature readings for each of the three thermocouples within the interaction region.
In order to calibrate the measured $1f/2f$ lock-in ratio in terms of absolute radians (effectively measuring the constant $\kappa$ in Eq. 9), both before and after each such set of runs we induced a common angular ‘uncrossing’ in our polarimeter by using a computer-controlled stepper motor and differential micrometer to effect a small, controlled polarization rotation of roughly 5 mRad in our second polarizer (analyzer). This calibration procedure was performed twice before and twice after each data set. Since we measure the ratio of amplitudes, the exact magnitude of the mechanical uncrossing is not important, only that it is the same for both lasers. The precision and reproducibility of this procedure was aided by the microradian-level control of our mechanical lever-based polarizer mount Lacy2024 . As discussed below, we assigned calibration errors based on both the reproducibility of repeated calibration steps and the variance of the before vs. after data-set calibration exercises.
We typically re-optimized the optical system after one data set. Over the course of several weeks, we changed a variety of experimental parameters such as the magnitude of the DC Faraday field, the E1 laser power, laser sweep conditions, etc. Data were taken over a wide range of sample temperatures between 700-$850\,^{\circ}$ which corresponds in both comparison experiments to an order-of-magnitude change in Pb vapor density, and a factor of four change in the thermal Boltzmann factor noted in the previous section. In the end, we collected and analyzed more than 500 individual spectral scan pairs for each comparison experiment taken over a wide variety of experiment conditions including a range of different DC magnetic fields and E1 laser powers. A comprehensive analysis of both statistical and potential systematic errors associated with the overall data sets is discussed below.
@@ IV.1 Line shape fitting procedure
Figure 4: Sample data (blue markers) for single (field on - field off) Faraday rotation scans for all three transitions at $\,{}^{\circ}$ and 44 Gauss. Each scan displayed required roughly 30 seconds to acquire. The line of best fit is shown in red, the expanded fit residuals are shown below each line shape (see text). While the 1279 nm and 368 nm optical rotation amplitudes are of comparable size, the amplitude of the 406 nm transition is an order of magnitude smaller due to the reduced thermal population of $(6p^{2})^{3}P_{2}$ level compared with the $(6p^{2})^{3}P_{1}$ level.
@@
Figure 5: Sample single transmission data scans (blue markers) and line of best fit (red) at $\,{}^{\circ}$ for the 368 nm (left) and 1279 nm (right) transition. The somewhat higher noise level for the 1279 nm transition signal is a consequence of the much smaller $2f$ modulation depth due to the reduced Verdet constant of the CeF3 crystal at 1279 nm.
@@
As a first step in preparing our spectra for analysis, a linearized and calibrated frequency axis of each spectra is obtained by fitting the Fabry-Pérot transmission data to an Airy distribution. Nonlinear scan behavior was accounted for by expanding the Airy argument up to a fourth-order polynomial. Higher-order polynomials were explored but were found to have no effect on the extracted E1 transition matrix element. Next, we subtract the lock-in $1f/2f$ signals with magnetic field off from the associated field-on spectra which removes frequency-dependent polarization backgrounds unrelated to the atoms. We then used the average value of the before- and after-data set calibration steps to provide background-subtracted spectra with amplitudes in units of radians. Using the calibrated frequency scale, we can separately create transmission spectra suitable for direct optical depth analysis using the $2f$ lock-in output for the case of zero magnetic field (that is, using our ‘background’ scans).
@@
We establish a nominal temperature to associate with each scan for two reasons. First, as described next, we are able to fix the Doppler width of our spectra simplifying the fitting process. Second, as shown in Eqs. 4 and 8, we require a temperature value to evaluate the Boltzmann factor that allows us to ultimately extract the E1 amplitude values. For the large majority of the data collected, the three thermocouple readings within the vapor cell enclosure were consistent within the quoted ±$\pm 1.5\,^{\circ}$ thermocouple accuracy. Here we assigned the central thermocouple value (labeled “B” in Fig. 3) to be the nominal temperature. We ultimately assigned an overall temperature-uncertainty systematic error to our results, as described in Sec. IV.C.
@@
Finally, we also establish a DC magnetic field value to associate with each Faraday scan so that we can fix the Zeeman splitting in our fits as well. In Appendix B, we describe our method to calibrate the magnetic field based on a Doppler-free, two-step spectroscopy scheme using the 1279 nm M1 laser overlapping with the 368 nm UV laser to probe the Zeeman structure of the $(6p^{2})^{3}P_{1}$ whose magnetic $g$-factor is very well known (see Wood1968 and references therein). Sec. IV.C. further discusses the accuracy of this method and potential systematic errors associated with possible inaccuracies in assigning magnetic field values to our fits.
@@
When fitting direct transmission spectra, we similarly fixed the Doppler widths and then fit the (magnetic field-off) $2f$ lock-in signals to the function described by Eqs. 1 and 2. We allowed for a frequency-dependent linear intensity background in the fits to account for incident laser power variation across the resonance. From these fits, we again extracted homogeneous width components and also peak optical depths from which the E1 amplitudes are obtained. Importantly, we saw good agreement of the homogeneous widths as extracted from the Faraday and the transmission line shapes.
For experimental conditions where direct transmission analysis was statistically feasible, these quite distinct analysis procedures provide a powerful test of potential hidden systematic errors. For example, the transmission fits do not rely on calibration of the signal amplitude or magnetic field, whereas the Faraday technique provides a zero-background signal whose very high intrinsic precision allows data collection over a much wider range of temperatures and vapor densities. For the 368 nm transition, originating from the $(6p^{2})^{3}P_{1}$ state, we were able to obtain high-quality transmission fits and extract accurate E1 amplitude values for the data taken at $\,{}^{\circ}$ and $\,{}^{\circ}$. Because of the order-of-magnitude smaller thermal Boltzmann factor for the $(6p^{2})^{3}P_{2}$ state, we were able to obtain transmission results with precision sufficient for useful statistical comparison to the Faraday results only at the highest temperature ($\,{}^{\circ}$) for 406 nm E1 transition.
Fig. 4 shows samples of single, background-subtracted Faraday rotation scans with solid lines representing model fits. These sample scans were taken with nominal values of magnetic field and laser power, and at a temperature in the middle of our experimental range. Each spectrum required roughly 30 seconds to acquire. Shown below the data and fit line are residuals, expanded by a factor of 10. The increased statistical noise in these residual plots align with the slopes of the Faraday curves, and is consistent with a frequency ‘jitter’ of 1-2 MHz, typical of our ECDL systems, as further discussed in Maser2019 . Fig. 5 shows a sample of direct transmission spectra for the 368 nm E1/M1 comparison experiment taken at $\,{}^{\circ}$. Data points and model fit are again shown.
@@ IV.2 Summary of results
After determining amplitude values from the Faraday fits (i.e. the $\mathcal{C}$ constant in Eq. 7), we insert the nominal temperature for the particular spectrum pair into the Boltzmann factor expression in Eq. 8 along with the other constants, and the accurately-known Porsev2016 M1 reduced matrix element $\langle||M1||\rangle=1.293(1)$ a.u. to obtain a value for the E1 reduced matrix element. When we analyze a full set of scans within one data set (typically 25 up and 25 down spectrum pairs), the fractional statistical standard error for the average extracted E1 transition amplitude is at or below $1\%$. Similarly, from a pair of fitted transmission spectra, we insert the values of the optical depths into Eq. 4 along with the thermal Boltzmann factor to again obtain E1 transition amplitudes.
@@
In Fig. 6 we have assembled results of all our data sets with associated error bars grouped in order of increasing cell temperature/density for each of the E1 transitions including Faraday results (red circles) as well the transmission line shape results (blue triangles). Over an order of magnitude change in Pb vapor density, we find no resolved trend in either E1 amplitude versus temperature, and that the transmission analysis is in excellent agreement with the Faraday results where common analysis was possible. Fig. 6 includes statistical errors only. For each data set we found the average (weighted by the statistical errors derived from each fit) and then added the associated fractional uncertainty of that set in quadrature with the uncertainty in our mechanical angle calibration procedure (typically a few tenths of 1% fractional uncertainty, and common to all scans within a set). For the latter contribution, we considered the difference between before-set and after-set calibration procedures.
As indicated in Table 1, the final fractional statistical error for each E1 transition amplitude was near 0.2%. We determined each final value by taking the weighted mean value of all data sets. As an alternative method, we measured the mean of the (unweighted) histogram consisting of every individual measurement of the transition amplitude. For both transitions, we found excellent agreement in the mean values derived from these two methods. Our final values (including statistical errors only) are: $\langle||E1||\rangle_{368}=1.897(4)$ a.u. and $\langle||E1||\rangle_{406}=3.008(7)$ a.u.
@@ IV.3 Systematic error discussion
We took two different approaches in exploring our results for potential systematic errors. First, we split the collected data into discrete subsets taken with differing experimental conditions and looked for evidence of statistically significant dependencies on various parameters. Examples of this include considering results from upward vs. downward laser scans, results for various values of DC magnetic fields used for Faraday rotation, different E1 laser powers, and, as plotted in Fig. 6, results grouped by cell temperature. Second, to explore possible line shape model or fitting inaccuracies, we refit subsets of data with altered model parameters or fitting procedures and looked for changes in the extracted values for our E1 matrix elements.
Although the large majority of the complete body of data were Faraday rotation-based, we nevertheless separated the analysis of the Faraday and transmission results to account for distinct systematic errors. While the Faraday results were far more statistically precise, they were also subject to more potential systematic errors. We here enumerate and summarize the systematic error contributions to the Faraday data only in Table 1. We note that the final two entries in the table (in red text) are factors which directly scale the extracted E1 amplitude (see Eqs. 4 and 8), and would potentially bias both Faraday and transmission data in the same direction. We therefore took the approach of analyzing statistical and all relevant systematic contributions for Faraday data and transmission data separately, took the weighted average of these, and then finally combined that resultant error with the contributions due to temperature and M1 amplitude uncertainty in quadrature to obtain the final quoted uncertainty discussed in Sec. V. Below we focus largely on the Faraday data systematics investigations in particular.
@@ IV.3.1 Line shape model and fits
Line shape fits were also performed both with and without the off-resonant ‘wings’ of the Faraday spectra, and while the observed changes were small, the effect on the 368 nm Faraday spectra was slightly larger, due to the complications associated with small ellipticity and asymmetry caused by the CeF3 crystal modulator Lacy2024 . For the 368 nm Faraday line shape fits, this ellipticity -- which arises from small differences in the amplitudes of the $\sigma_{+}$ and $\sigma_{-}$ contributions to the (otherwise symmetric) Faraday spectra -- was explicitly accounted for by the introduction of a small ellipticity parameter in the model. Typically, these effects were statistically unresolved, which was achieved by ensuring the lasers were aligned along the optic-c axis of the crystal (see Lacy2024 for further details). To explore possible correlations of this ellipticity with Faraday amplitude and eventually matrix element determination, we compared fits of a representative batch at $\,{}^{\circ}$ both with and without an ellipticity fit parameter and found changes at the 0.1% level. Accounting for such ellipticity was not required for the 406/1279 nm $\langle\norm{E1}\rangle$ comparison, since the biaxial TGG crystal modulator used there exhibits no birefringenece.
Another potential feature in all of our Faraday line shapes is due to possible imperfect background subtraction which we explored by re-fitting subsets of data with a linear background slope parameter. The results of these studies are included in Table 1. We note that the first three entries in the table apply also to the analysis of possible systematic errors in our transmission spectra, and we took a similar approach to re-fitting these spectra, with error contributions at a similar 0.1% level or below.
@@ IV.3.2 Calibration uncertainties
In our Faraday rotation line shape analysis, we enter the particular Zeeman shifts as fixed, ‘known’ parameters while we extract the line shape amplitude and homogeneous widths from the fits themselves. If either the magnetic field calibration or the values of the relevant $g$-factors are incorrect, this could lead to systematic errors in key fitted parameters. The various $g$-factors discussed and referred to in Sec II have been measured Wood1968 ; Lurio1970 and are in good agreement with theory values Porsev2016 . We use the following consensus values: $g=1.501(2);\,g^{\prime}=1.276(2);\,g^{\prime\prime}=1.352(3)$. Residual errors in these values do not contribute to the 368/1279 nm comparison since they involve a common $g$-factor. To study the impact of these uncertainties on the 406/1279 nm comparison, using both experimental and simulated data, we refit the spectra with $g$-factors altered by an amount consistent with the error bars stated above, and assessed the impact of these uncertainties on the eventual computed value for the E1 amplitude (see Table 1).
Regarding magnetic field calibration, our in situ method for field calibration at the 1% level of accuracy is described in Appendix B. We can similarly explore the consequence of incorrectly fixed magnetic field values through refitting experimental or simulated data with various slightly altered fixed values for the field. In the limit of the small-field approximation, the effect of a changing magnetic field would exactly cancel in our ratio measurement; here there is some differential effect since the Zeeman splitting represents a different fraction of the overall line shape width for the M1 vs E1 cases.
@@
The most significant uncertainty in our entire experimental project, equally relevant for Faraday and Transmission data analysis, is associated with temperature calibration and the effect that this has on the thermal Boltzmann factor appearing in Eqs. 4 and 8. As discussed in Sec. III, the sample temperature was monitored with three s-type thermocouples located in the stainless stell enclosure containing our vapor. These have a quoted accuracy of $\Delta T=\pm$ $\,{}^{\circ}$. In most cases, the three thermocouple readings were in agreement within this intrinsic error. In these instances, we take the temperature uncertainty as $\Delta T/\sqrt{3}\approx\pm 0.9$$\,{}^{\circ}$. For the 368 nm and 406 nm transitions, the corresponding uncertainty in the relevant Boltzmann factors leads to an $\langle\norm{E1}\rangle$ uncertainty of 0.46% and 0.58% respectively at $\,{}^{\circ}$ where the bulk of the data were collected (this percentage error depends very weakly on temperature over our full experimental range). In the few cases where there was a small but resolvable temperature gradient across the atomic sample, we assigned an additional temperature uncertainty (up to $\pm$ $\,{}^{\circ}$) whose additional propagated uncertainty in the transition amplitude value was added in quadrature with fit errors and calibration errors, effectively reducing the statistical weight of that run. This additional gradient uncertainty was estimated by finding the difference between the central thermocouple reading, $T_{B}$, and the integrated average of the temperature distribution across the length of the cell. The temperature distribution was found by fitting the three thermocouple readings, $T_{A,B,C}$, considering their locations with respect to the cell (see Fig. 3), and assuming a smooth, continuous function of position.